OpenStudy (amonoconnor):

When using the Chain Rule, and taking the second piece of derivative (dy/dx of the term for "theta"), would this be correct... (?) In my particular situation: The "theta" = (1/y) So... Because it's not the first step in the chain rule, and I'm "re-deriving" as my math teacher says, the constant (1) is affectable. So... Is the following derivation correct? 1/y = (0)/[y^(1-1)(dy/dx)] = 0/y^(0)(dy/dx) = 0/(1)(dy/dx) = 0/(dy/dx) = (0)

OpenStudy (amonoconnor):

@agent0smith Yes... I know. I'm a thorn in your side ;)

zepdrix (zepdrix):

"theta" = (1/y) ?? I don't understand :(

OpenStudy (amonoconnor):

Yep, in he function I'm taking the derivative of, the quantity "1/y" is in the "inside" spot, in Chain Rule terminology, or the spot of theta in regard to the trig function also in the function. I have taken the derivative of the whole "outside", finding the derived/ new trig function (theta term untouched), but now I must take the derivative of the theta term itself, and multiply that by the first part; to complete the derivation process, and find the final answer. I'm just wondering if this is in fact the right derivative. :)

zepdrix (zepdrix):

Ok so for the derivative of the inner function (1/y), You're thinking you get zero for some reason? Did you apply the quotient rule or power rule or something?

OpenStudy (amonoconnor):

My problem: "ySIN(1/y) = 1 - xy"; Find dy/dx. ( Thought this would be helpful to see)

OpenStudy (amonoconnor):

Umm, no.... I should have, shouldn't I have though? Used the quotient rule?

zepdrix (zepdrix):

You can rewrite it like this:\[\Large \frac{1}{y}\quad=\quad y^{-1}\]And apply the power rule as you normally would. I'm not quite sure where your 0 is coming up in the denominator. If you use the quotient rule you should get something like:\[\Large \frac{0y-1(1)}{y^2}\]

zepdrix (zepdrix):

the zero in the numerator, i meant to say :)

zepdrix (zepdrix):

Oh and a \(\Large y'\) should show up when you chain again.

zepdrix (zepdrix):

But it looks like you understood that part :o i see your dy/dx in there somewhere (although it shouldn't be in the denominator XD heh )

OpenStudy (amonoconnor):

The "Power Rule"? I'm not familiar with that... This is what we learned in Calc, at least from my teacher: (?) y' = (y^(1-1))*(dy/dx) = (y^(0))*(dy/dx) = (1)*(dy/dx) = dy/dx

zepdrix (zepdrix):

Is that a y to the first power on the left there?

OpenStudy (amonoconnor):

Y prime

zepdrix (zepdrix):

You've learned about the chain rule before the power rule...?\[\Large \frac{d}{dx}x^n\quad=\quad n x^{n-1}\]That doesn't look familiar? +_+

OpenStudy (amonoconnor):

Oh... Yes, I know of that. Whoops! I guess we don't know the terminology... I will bring that up to him!

OpenStudy (amonoconnor):

Ahhhh, I follow your post from earlier... :)

zepdrix (zepdrix):

Using rules of exponents allows us to write the inner function like this:\[\Large \frac{1}{y}\quad =\quad y^{-1}\]From here we don't need to use the quotient rule, we use the power rule instead:\[\Large \frac{d}{dx}y^{-1}\quad=\quad -1y^{-1-1}\frac{dy}{dx}\quad=\quad -y^{-2}\frac{dy}{dx}\]

OpenStudy (amonoconnor):

Alright.... You're so amazing. I do believe I'm good now. I have what I need, and most importantly, I fully understand what happened, and what I need to do. Thank you so much!

OpenStudy (amonoconnor):

(Would you mind medaling me?)

zepdrix (zepdrix):

oh got it from there? cool c:

OpenStudy (amonoconnor):

Yep, I do! Thanks you so much again; you really helped me look at that from a better angle.